# Turning the Tide on Dark Pools? Towards Multi-Stakeholder Vulnerability Notifications in the Ad-Tech Supply Chain

### v1

@@ 4 Recipient Engagement with Notifications
Our study included a total of 2.9K unique entities who were involved in or impacted by dark pooling in the online advertising supply chain. These included 1.7K victim publishers, 644 ad-networks, and 635 advertisers. In this section, we report the rates of engagement (Section 4.1) and the nature of engagement (Section 4.2) with our notifications by these entities.
@@ 4.1 Rate of engagement with notifications
Notification delivery rates. Of the 2.9K entities involved in our study, we only sent notifications to two-thirds (cf. Section 3.2) and used the remaining one-third as a control group against which the effect of our notifications could be determined. Of these 1.9K notifications (across both treatment groups $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$), 79.5% were successfully delivered and opened by their recipients. This rate of delivery exceeds the rates reported in prior notification research [28, 27, 24, 38]. We hypothesize that this is because of our ability to leverage email addresses from ads.txt and sellers.json files for contacting publishers and ad-networks, respectively. The remaining notifications (417 notifications or 20.5%) failed to be delivered for a variety of reasons including emails rejected by spam filters (40.2% of failures) and invalid email address (21.5% of failures).
Notification response rates. Of the 1.5K recipients who read our notifications, we received an email response from 215 (14%). Of these, 147 emails were generic responses that appeared to be template responses. The remaining 68 emails were engaging directly with the content of our notifications and were qualitatively analyzed (cf. Section 4.2). Of these, 40 were from victim publishers (3.5% response rate), 18 were from ad-networks (4.2% response rate), and 10 were from advertisers (2.9% response rate). This is in line with the response rates observed in prior research [30, 31, 38, 44, 45, 46] focused on single-stakeholder notifications. It should be noted that the absence of an email response does not suggest that the notifications did not influence vulnerability resolution. We measure the exact influence on vulnerability resolution in Section 5.
@@ 4.2 Nature of engagement with notifications
Methodology for thematic analysis. To systematically explore different themes emerging in responses to our notifications, we adopted the state-of-the-art thematic analysis methodology [62, 44] to identify, analyze, and report patterns in the responses. This analysis can uncover common attitudes towards and challenges faced in the ad-tech vulnerability resolution. One author familiarized themselves with the email responses by reading them multiple times. Next, a codebook was developed by the same author by assigning codes to each sentence in the response emails. This codebook was iteratively refined, in consultation with the other authors, to consolidate the final set of codes in our codebook. Next, the authors use this codebook to independently code all the emails. Following Clarke and Braun’s [63] understanding of thematic analysis, we also did not calculate inter-rater reliability (similar to Stover et. al [44], which used the same thematic analysis methodology). Instead, the authors convened to deliberate any ambiguities and reach a consensus on the final coding in order to ensure a high quality of analysis. Finally, the email responses were grouped into three high level themes: (1) effort towards resolution, (2) demonstration of concern or awareness, and (3) lack of trust or resources. The codebook, identified themes, categories, and sub-categories is presented in Table II. Next, we use the developed codebook categories to study responses by themes, recipient type, and sender reputation (Cf. Figure 5 in the appendix).
TABLE II: Code book with themes, categories, and subcategories that resulted from the thematic analysis of notification responses.
Category Subcategories Frequency
@@
Fixed reported issues made fixes (7); asked to verify current status of issues post resolution (2); investigated the issues (2); removed off the vulnerable entries present in the report (2); actively fixing the issue (1); ads.txt added but never used (1); sellers.json had stale entries (1); will not engage in further correspondence on security issues (1) 11
@@
Justifying the issue defending few reseller relations (1); justifying low inventory fraud rate due to the usage of brand-safety tools (1); questioning vulnerabilities as benign due to direct relations (1); stating to work via mostly direct-business relations (1) 4
@@
Motivated to fix will investigate on the issue (13); will resolve the issue (8); willing to help fix the issue (3); will remove ads.txt (1); will remove vulnerable entries from sellers.json (1) 21
@@
Response theme 1: Effort towards resolution. We found that victim publishers and ad-networks often requested additional actionable information or clarifications to aid the resolution of our reported issues. This is not commonly seen in responses from advertisers, however. We hypothesize that this is because advertisers are less likely to maintain in-house tech-teams and instead rely on ad agencies to manage their ad campaigns for them. This hypothesis is supported by the high number of forwarded to the right department category responses observed from advertisers. This suggests that most of the generic contact emails available on a advertiser’s website may not be particularly useful for performing advertiser notifications related to ad-fraud. In contrast, victim publishers yielded a mix of responses under this category – smaller publishers did not have separate teams for advertising, however, larger publishers managed advertising under a separate team to which our notifications were forwarded. We found that many of the responsive victim publishers and ad-networks were motivated to collaboratively resolve the issues and scheduled either a video meeting or a phone call to understand and resolve the reported vulnerability. Victim publishers were most likely to have responded following a successful resolution of the reported vulnerability. For example, one victim publisher responded – “We’ve cleaned up our ads.txt to only active providers. Can you check it again?” – suggesting that they do not regularly update their ads.txt files to maintain them up-to-date. Untimely updates can aid problematic publishers to monetize through victim publisher’s sellerIDs.
@@
Response theme 3: Lack of trust or resources. Several entities highlighted their inability or unwillingness to resolve the reported vulnerabilities for a variety of reasons ranging from doubting the correctness of the report to a lack of technical resources to properly understand and resolve the vulnerabilities. In one interesting case, an ad-network claimed our report (which contained evidence that a misinformation publisher was pooling a seller ID issued by them) was inaccurate — “My tech team verified that all the references on your list are legitimate publishers owned by one of the corporate entities on our small inclusion list”; yet, within hours, this same ad-network updated their sellers.json file with a comment “We no longer work with [misinformation publisher] or any of its associated properties.” Several entities also suspected that our notifications were part of a phishing campaign and accordingly delayed resolution actions. These entities often sought proof of our identities, with several even reaching out to our university’s media relations team for clarification. “… a lawyer who received an email from your address yesterday is concerned about the issue that this email may have been a phishing attempt.”. Fortunately, these suspicions were eventually resolved favorably. Finally, many publishers reported their inability to resolve the reported vulnerabilities due to the lack of technical resources — for example, one publisher reported “I’ve made some immediate fixes — basically, just cleaning up our ads.txt file, removing the lines in question but also deleting a lot of probably unnecessary other lines as well. Beyond that, I’m not sure how to contact all the various other players — the ad-networks, the problematic websites. We’re a very a small publisher…”. This lack of technical know-how was also found to be the cause of several reported vulnerabilities, as one publisher reported – “I had no idea that there was any issue with our ads.txt file. I’ve always just included whatever lines that [AdX] tell us to add.
@@ 5 Notification Impact on Dark Pooling
Our qualitative analysis on responses to notifications does not provide a complete picture on their impact because recipients may initiate remedial actions without responding to the notification emails. In this section, we examine two metrics for assessing the influence of notifications (Section 5.1). We then use these to measure how the notification source (Section 5.3) and recipient (Section 5.2) influence remedial responses.
@@ 5.1 Assessing the impact of notifications
Figure 4: An illustration of the measured remediation metrics before and after notifications in different rounds to different stakeholders. The difference-in-differences is computed from these metrics as $\Delta_{(t,c)}=(t_{post}-c_{post})-(t_{pre}-c_{pre})$ for each ad-tech entity.
Propensity matched difference-in-differences (PM-DiD). Our goal is to uncover the overall impact of notifications on specific measures of dark pooling relevant to each stakeholder. We use propensity-matched difference-in-differences, a common statistical approach for identifying the causal effect size of an intervention when a randomized controlled trial is not possible [64, 65, 66]. We apply PM-DiD in 3 steps.
Step 1: Matching entities in treatment and control groups. First, we create a control group that is statistically similar to the treatment group based on specific dark pooling measures. Our treatment group may be $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, or $\mathcal{T}_{1}\bigcup\mathcal{T}_{2}$ and our control group is some subset of $\mathcal{C}$ (cf. Section 3.2). For each entity in our treatment group, we find the entity in the control group candidates which is the most similar for the given measures (i.e., its pre-notification nearest neighbor) and include this candidate in our control group.
Step 2: Computing pre- and post-intervention measures. Next, we identify the pre- and post-intervention measures for each entity in the treatment group ($t$) and their matched control counterpart ($c$). We record pre-intervention measures ($t_{pre}$ and $c_{pre}$) from crawls that occur 1-2 weeks before sending notifications to the treatment entity. We record post-intervention measures ($t_{post}$ and $c_{post}$) from crawls that occur 4-5 weeks after sending notifications. An illustration of effective remediations is provided in Figure 4.
@@
Measures of dark pooling for each stakeholder. We track the following dark pooling metrics to measure the impact of our notifications on each stakeholder.
Victim publishers. We focus on one key metric: the number of problematic publishers which pool and use the seller ID of the victim publishers (we refer to this count as probdomainsvp for a publisher $vp$). We expect probdomainsvp to reduce if $vp$ engages with their ad-networks to either get: (1) an unpooled seller ID, (2) assigned a different pool than the problematic publishers, or (3) remove problematic publishers from their pool.
Ad-networks. We use two metrics for ad-networks: (1) the number of seller IDs they issue that are used for dark pools (we refer to this as poolsad for the ad-network $adx$); and (2) the number of seller IDs issued by other ad-networks to $adx$ that facilitate dark pooling (we refer to this as partnerpoolsad for $adx$). We expect poolsad to decrease if $adx$ corrects its own pool assignments to prevent pooling of problematic and popular domains. We expect partnerpoolsad to decrease if $adx$ engages with other ad-networks where it is listed to correct their facilitation of dark pools.
Advertisers. Our notifications to advertisers only informs them about the ad-network responsible for displaying their ads on problematic publishers and the associated seller ID. Therefore, we expect a reduction in poolsad for the responsible $adx$ if the advertisers request to remove such poolsad. The fact that advertisers monetize the ad-tech supply chain would motivate ad-networks to act.
@@ 5.2 Effect of notification recipient
Recipient (Metric) $N_{\text{rem.}}$ $\mu_{\text{ov.}}$ $\mu_{\text{rem.}}$
@@
TABLE III: Effect of notification recipient on dark pooling.
Table III shows the impact of notifications on dark pooling metrics for each stakeholder. The data reveals several interesting findings. Notably, all stakeholders are generally responsive to our notifications with between 54-82% of them showing a remedial action (i.e., $\Delta_{(t,c)}<0$).
Ad-networks are most responsive; publishers are least responsive. At the less responsive end, 54% of notified publishers took successful remedial action, reducing their participation in dark pools more than their control counterparts. In contrast, 81.6% of notified ad-networks took successful remedial action, reducing the number of seller IDs used for dark pooling at a higher rate than their control counterparts. Rather surprisingly, despite controlling the revenue of the advertising supply chain, only 72% of advertisers showed some mitigation in dark pools. This finding might be explained by the fact that, among all the stakeholders, ad-networks can mitigate dark pools most easily because they control their own assignment of seller IDs. On the other hand, publishers and advertisers need to engage with their ad-networks. Our qualitative analysis supports this, showing that many publishers were unsure how to resolve their participation in dark pools. Moreover, this analysis shows that ad-networks remediate dark pools with a higher probability when advertisers reach out to them as compared to when publishers do so – likely because advertisers are actually the source of funding in the ad-tech supply chain.
@@ 5.3 Effect of notification source
Source Recipient (Metric) $N_{\text{rem.}}$ $\mu_{\text{ov.}}$ $\mu_{\text{rem.}}$
@@
TABLE IV: Effect of notification source on measures of dark pooling for each stakeholder. denotes a statistically significant ($t$-test; $p<$.05) difference in the overall effect based on notification source.
@@ 6 Discussion and Conclusion
In this section, we consider the validity of our study design (Section 6.1), the limitations of our research (Section 6.2), ethical considerations (Section 6.3), and implications for multi-stakeholder notification studies (Section 6.4)
@@ 6.1 Validity of study design
Internal and external validity considerations. Internal validity measures how well a study is conducted and how accurately its results reflect the studied group, while external validity assesses the generalizability of the findings. We evaluate different internal and external validity threats applicable to our study and explain how we handled them to minimize any validity issues. We follow recommendations from prior works [67, 68] to ensure the validity of our study and results.
Threats from lack of representativeness. We ensured representativeness by notifying all entities found to participate in dark pooling, without imposing selection criteria for ad-networks and advertisers. However, problematic publishers often impersonate popular publishers. So, we selected our sample from the Tranco top-10K publishers whose advertising identifiers were used by problematic publishers in their ads.txt files. We randomly assigned the selected entities into different treatment and control groups to avoid selection bias. We applied the same treatment steps to different entities, though the specific vulnerabilities disclosed and remediation steps suggested varied based on their role in the ad-tech supply chain. We did not introduce any other experimenter bias.
Threats from subject attrition. We did not measure the long-term effects of notifications on entities, so attrition concerns are not relevant. Additionally, ad-tech entities do not change roles over time (e.g., a publisher does not become an ad-network). We ensured blinding by not informing participants about the applied treatment or intervention. We did not perform any experimental manipulation, such as entitiy-specific priming or incentives.
Threats from situational factors. Our notified entities are spread across the globe. We did not account for situational factors like the time and date of notification or location, which could affect the external validity of our results. We acknowledge this limitation.
Threats from lack of trust and realism. To improve external validity, we aimed for psychological realism [69] by communicating through our email text and website content that we were working in the entity’s favor, aiming to reduce their participation in a problematic activity. This approach developed their trust in us, increasing confidence in our notifications. Our notifications were also based on verifiable real data which reflected their true participation in dark pools, further bolstering realism and building trust.

### v2

@@ 4 Recipient Engagement with Notifications
Our study included a total of 2.9K unique entities involved in or impacted by dark pooling in the ad-tech supply chain. These comprise of 1.7K victim publishers, 644 ad-networks, and 635 advertisers. In this section, we report the rates (Section 4.1) and the nature (Section 4.2) of notification engagement.
@@ 4.1 Rate of engagement with notifications
Notification delivery rates. Of the 2.9K entities involved in our study, we only sent notifications to two-thirds (cf. Section 3.2) and used the remaining one-third as a control group against which the effect of our notifications could be determined. Of these 1.9K notifications (across both treatment groups $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$), 79.5% were successfully delivered and opened by their recipients. This rate of delivery exceeds the rates reported in prior notification research [cetin2017make, cetin2016understanding, stock2016hey, stock2018didn]. We hypothesize that this is because of our ability to leverage email addresses from ads.txt and sellers.json files for contacting publishers and ad-networks, respectively. The remaining notifications (417 notifications or 20.5%) failed to be delivered for a variety of reasons including emails rejected by spam filters (40.2% of failures) and invalid email address (21.5% of failures).
Notification response rates. Of the 1.5K recipients who read our notifications, we received an email response from 215 (14%). Of these, 147 emails were generic responses that appeared to be template responses. The remaining 68 emails were engaging directly with the content of our notifications and were qualitatively analyzed (cf. Section 4.2). Of these, 40 were from victim publishers (3.5% response rate), 18 were from ad-networks (4.2% response rate), and 10 were from advertisers (2.9% response rate). This is in line with the response rates observed in prior research [maass2021snail, maass2021effective, stock2018didn, stover2023website, soussi2020feasibility, utz2023comparing] focused on single-stakeholder notifications. It should be noted that the absence of an email response does not suggest that the notifications did not influence vulnerability resolution. We measure the exact influence on vulnerability resolution in Section 5.
@@ 4.2 Nature of engagement with notifications
Methodology for thematic analysis. To systematically explore different themes emerging in responses to our notifications, we adopted the state-of-the-art thematic analysis methodology [braun2006using, stover2023website] to identify, analyze, and report patterns in the responses. This analysis can uncover common attitudes towards and challenges faced in the ad-tech vulnerability resolution. This approach eliminates the need for multiple coders, so one author familiarized themselves with the email responses by reading them multiple times. Next, a codebook was developed by the same author by assigning codes to each sentence in the response emails. This codebook was iteratively refined, in consultation with the other authors, to consolidate the final set of codes in our codebook. This is because [clarke2013successful] emphasizes that thematic analysis is reflexive and interpretive rather than a reliability exercise, so one researcher can reasonably undertake initial coding to ensure immersion and consistency, with credibility and analytic quality ensured through subsequent collaborative refinement and consensus, eliminating the need for statistical reliability. Next, the authors use this codebook to independently code all the emails. We follow Stover et. al [stover2023website], who adopt Clarke and Braun’s [clarke2013successful] understanding of thematic analysis without requiring to calculate inter-rater reliability. Instead, similar to their work, the coder convened with the second author to deliberate any ambiguities and reach a consensus on the final coding in order to ensure a high quality of analysis as suggested in the approach. Finally, the email responses were grouped into three high level themes: (1) effort towards resolution, (2) demonstration of concern or awareness, and (3) lack of trust or resources. The codebook, identified themes, categories, and sub-categories are presented in Table II. Next, we use the developed codebook categories to study responses by themes, recipient type, and sender reputation (Cf. Table V (appendix)).
TABLE II: Code book with themes, categories, and subcategories from the thematic analysis of notification responses.
Category Sub-categories Frequency
@@
Fixed the reported issues made fixes (7); asked to verify current status of issues post resolution (2); investigated the issues (2); removed off the vulnerable entries present in the report (2); actively fixing the issue (1); ads.txt added but never used (1); sellers.json had stale entries (1); will not engage in further correspondence on security issues (1) 11
@@
Justifying the reported issue defending few reseller relations (1); justifying low inventory fraud rate due to the usage of brand-safety tools (1); questioning vulnerabilities as benign due to direct relations (1); stating to work via mostly direct-business relations (1) 4
@@
Motivated to fix the issue will investigate on the issue (13); will resolve the issue (8); willing to help fix the issue (3); will remove ads.txt (1); will remove vulnerable entries from sellers.json (1) 21
@@
Response theme 1: Effort towards resolution. We found that victim publishers and ad-networks often requested additional actionable information or clarifications to aid the resolution of our reported issues. This is not commonly seen in responses from advertisers, however. We hypothesize that this is because advertisers are less likely to maintain in-house tech-teams and instead rely on ad agencies to manage their ad campaigns for them. This hypothesis is supported by the high number of forwarded to the right department category responses observed from advertisers. This suggests that some of the generic contact emails from an advertiser’s website may not be particularly useful for performing advertiser notifications related to ad-fraud. In contrast, victim publishers yielded a mix of responses under this category – smaller publishers did not have separate teams for advertising, however, larger publishers managed advertising under a separate team to which our notifications were forwarded. We found that many of the responsive victim publishers and ad-networks were motivated to collaboratively resolve the issues and scheduled either a video meeting or a phone call to understand and resolve the reported vulnerability. Victim publishers were most likely to have responded following a successful resolution of the reported vulnerability. For example, one victim publisher responded – “We’ve cleaned up our ads.txt to only active providers. Can you check it again?” – suggesting that they do not regularly update their ads.txt files to maintain them up-to-date. Untimely updates can aid problematic publishers to monetize through victim publisher’s sellerIDs.
@@
Response theme 3: Lack of trust or resources. Several entities highlighted their inability or unwillingness to resolve the reported vulnerabilities for a variety of reasons ranging from doubting the correctness of the report to lack of technical resources to properly understand and resolve the vulnerabilities. In one interesting case, an ad-network claimed our report (which contained evidence that a misinformation publisher was pooling a seller ID issued by them) was inaccurate — “My tech team verified that all the references on your list are legitimate publishers owned by one of the corporate entities on our small inclusion list”; yet, within hours, this same ad-network updated their sellers.json file with a comment “We no longer work with [misinformation publisher] or associated properties.” Several entities also suspected our notifications to be part of a phishing campaign. These entities often sought proof of our identities, with some even reaching out to our university’s media relations team for clarification. “… a lawyer who received an email from your address yesterday is concerned about the issue that this email may have been a phishing attempt.”. Fortunately, these suspicions were eventually resolved favorably. Finally, many publishers reported their inability to resolve the reported vulnerabilities due to the lack of technical resources — for example, one publisher reported “I’ve made some immediate fixes — basically, just cleaning up our ads.txt file, removing the lines in question but also deleting a lot of probably unnecessary other lines as well. Beyond that, I’m not sure how to contact all the various other players — the ad-networks, the problematic websites. We’re a very a small publisher…”. This lack of technical know-how was also found to be the cause of several reported vulnerabilities, as one publisher reported – “I had no idea that there was any issue with our ads.txt file. I’ve always just included whatever lines that [AdX] tell us to add.
@@ 5 Notification Impact on Dark Pooling
Our qualitative analysis on responses to notifications does not provide a complete picture on their impact because recipients may initiate remedial actions without responding to the notification. In this section, we examine 3 measures for assessing the influence of notifications (Section 5.1). We then use these to evaluate how the notification source (Section 5.3) and recipient (Section 5.2) influence remedial responses.
@@ 5.1 Assessing the impact of notifications
Figure 4: An illustration of the measured remediation measures before and after notifications to different stakeholders in different rounds. The difference-in-differences is computed from these measures as $\Delta_{(t,c)}=(t_{post}-c_{post})-(t_{pre}-c_{pre})$ for each ad-tech entity.
Propensity matched difference-in-differences (PM-DiD). Our goal is to uncover the overall impact of notifications on specific measures of dark pooling relevant to each stakeholder. We use propensity-matched difference-in-differences, a common statistical approach for identifying the causal effect size of an intervention when a randomized controlled trial is not possible [callaway2021difference, donald2007inference, tanveer2023glowing]. We randomly divide the notification recipients in each round into three equal-sized groups: two treatments and one control. We apply PM-DiD in 3 steps as shown in Algorithm 1 (Appendix D).
Step 1: Matching entities in treatment and control groups. First, we create a control group that is statistically similar to the treatment group based on specific dark pooling measures. Our treatment group may be $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, or $\mathcal{T}_{1}\bigcup\mathcal{T}_{2}$ and our control group is some subset of $\mathcal{C}$ (cf. Section 3.2). For each entity in our treatment group, and for each comparison measure, we find its pre-notification nearest neighbor—the entity in the control group candidates which is the most similar for the given measure (i.e., has the value of that measure numerically most close to the treatment entity in consideration) and include this candidate in our control group. This helps us to establish a baseline where the pre-notification differences between treatment and selected control are minimal and the effect observed post notification can be attributed to the performed treatment.
Step 2: Computing pre- and post-intervention measures. Next, we identify the pre- and post-intervention measures for each entity in the treatment group ($t$) and their matched control counterpart ($c$). We record pre-intervention measure ($t_{pre}$ and $c_{pre}$) from crawls that occur 1-2 weeks before sending notifications to the treatment entity. We record post-intervention measures ($t_{post}$ and $c_{post}$) from crawls that occur 4-5 weeks after sending notifications. Figure 4 shows an illustration of effective remediations.
@@
Measures of dark pooling for each stakeholder. We track the following dark pooling measures to evaluate the impact of our notifications on each stakeholder.
Victim publishers. We focus on one key measure: the number of problematic publishers which pool and use the seller ID of the victim publishers (we refer to this count as probdomainsvp for a publisher $vp$). We expect probdomainsvp to reduce if $vp$ engages with their ad-networks to either get: (1) an unpooled seller ID, (2) assigned a different pool than the problematic publishers, or (3) remove problematic publishers from their pool.
Ad-networks. We use two measures for ad-networks: (1) the number of seller IDs they issue that are used for dark pools (we refer to this as poolsad for the ad-network adxadx); and (2) the number of seller IDs issued by other ad-networks to adxadx that facilitate dark pooling (we refer to this as partnerpoolsad for adxadx). We expect poolsad to decrease if adxadx corrects its own pool assignments to prevent pooling of problematic and popular domains. We expect partnerpoolsad to decrease if adxadx engages with other ad-networks where it is listed to correct their facilitation of dark pools.
Advertisers. Our notifications to advertisers only informs them about the ad-network responsible for displaying their ads on problematic publishers and the associated seller ID. Therefore, we expect a reduction in poolsad for the responsible adxadx if the advertisers request to remove such poolsad. The fact that advertisers monetize the ad-tech supply chain would motivate ad-networks to act.
@@ 5.2 Effect of notification recipient
TABLE III: Effect of notification recipient on dark pooling.
Recipient (Measure) $N_{\text{rem.}}$ $\mu_{\text{ov.}}$ $\mu_{\text{rem.}}$
@@
Table III shows the impact of notifications on dark pooling measures for each stakeholder. The data reveals several interesting findings. Notably, all stakeholders are generally responsive to our notifications with between 54-82% of them showing a remedial action (i.e., $\Delta_{(t,c)}<0$).
Ad-networks are most responsive; publishers are least responsive. At the less responsive end, 54% of notified publishers took successful remedial action, reducing their participation in dark pools more than their control counterparts. In contrast, 81.6% of notified ad-networks took successful remedial action, reducing the number of seller IDs used for dark pooling at a higher rate than their control counterparts. Rather surprisingly, despite controlling the revenue of the advertising supply chain, only 72% of advertisers showed some mitigation in dark pools. This finding might be explained by the fact that, among all the stakeholders, ad-networks can mitigate dark pools most easily because they control the assignment of seller IDs. Ad-exchanges are the central entity in the ad-tech supply chain as they work with both sellers (i.e., publishers to auction their ad inventory) and buyers (i.e., advertisers to bid and buy relevant ad inventory from). Since dark pools are a result of shared usage of these seller IDs assigned by the ad-exchanges to different sellers, they can be reasoned to be the most responsive. On the other hand, publishers and advertisers need to engage with their ad-networks. Our qualitative analysis supports this, showing that many publishers were unsure how to resolve their participation in dark pools. Moreover, this analysis shows that ad-networks remediate dark pools with a higher probability when advertisers reach out to them as compared to when publishers do so – likely because advertisers are the source of funding in the ad supply chain.
@@ 5.3 Effect of notification source
TABLE IV: Effect of notification source on measures of dark pooling for each stakeholder. denotes a statistically significant ($t$-test; p<p<.05) difference in the overall effect based on notification source.
Source Recipient (Measure) $N_{\text{rem.}}$ $\mu_{\text{ov.}}$ $\mu_{\text{rem.}}$
@@ 6.1 Validity of study design
Internal and external validity considerations. Internal validity measures how well a study is conducted and how accurately its results reflect the studied group, while external validity assesses the generalizability of the findings. We evaluate different internal and external validity threats applicable to our study and explain how we handled them to minimize any validity issues. We follow recommendations from prior works [campbell2015experimental, anderson2018design] to ensure the validity of our study and results.
Threats from lack of representativeness. We ensured representativeness by notifying all entities participating in dark pooling, without imposing selection criteria for ad-networks and advertisers. But, problematic publishers often impersonate popular publishers. So, we selected our sample from the top-10K publishers whose advertising identifiers were used by problematic publishers in their ads.txt files. We randomly assigned the selected entities into different treatment and control groups to avoid selection bias. We applied identical treatment steps to different entities, though the specific vulnerabilities disclosed and remediation steps suggested varied based on their role in the ad-tech supply chain. We did not introduce any other experimenter bias.
Threats from subject attrition. We did not measure the long-term effects of notifications on entities, so attrition concerns are not relevant. Additionally, ad-tech entities do not change roles over time (e.g., a publisher does not become an ad-network). We ensured blinding by not informing participants about the applied treatment or intervention. We did not perform any experimental manipulation, such as entity-specific priming or incentives.
Threats from situational factors. Notified entities are globally spread. We did not account for situational factors like the time and date of notification or location, which could affect the external validity of our results. It’s a limitation.
Threats from lack of trust and realism. To improve external validity, we aimed for psychological realism [brewer2000research] by communicating through our email text and website content that we were working in the entity’s favor, helping to reduce their participation in a problematic activity. This approach was intended to develop their trust in us, increasing confidence in our notifications. Our notifications were also based on verifiable real data which reflected their true participation in dark pools, further bolstering realism and building trust.